1. Millicent is planning to travel to the United States of America in 3 years time. She 

estimated that her vacation would cost $1,236,000. Given the existing interest rate of 6%, 

how much money should she invest now?

A. $999,769.43

B. $1,027,869.53

C. $1,037,769.43

D. $1,250,869.53


1
Expert's answer
2022-05-16T18:55:01-0400

Solution

Let's use the formula to determine the future value of compound interest:

FV = PV*(1+r)^n

where FV is the future value;

       PV is present value;

       r is interest rate.

In our case, the FV is $1,236,000,  r is 6%, n is 3 years, and we need to find the present value (PV)

PV = FV / (1+r)^n = 1,236,000 / (1+0,06)^3 = $1,037,769.43

Millicent must invest $1,037,769.43 now to have $1,230,000 in 3 years at 6% per annum

Answer:C. $1,037,769.43